Could Use Help On Question. Thank You.Famas Llamas has a weighted average cost of capital of 9.6 percent. The companys cost of equity is 11 percent, and its pretCo
Answer #1
After tax rate = YTM * (1-Tax rate)
After tax rate = 6.9 * (1-0.24)
After tax rate = 5.24
Cost of Capital = Weight of Equity*Cost of Equity+Weight of Debt*Cost of Debt
9.6 = 11*Weight of Equity+5.24*(1-weight of Equity)
Weight of Equity = 0.7569
Weight of Debt =1-weight of Equity=1-0.7569=0.2431

D/E = 0.2431/0.7569

=0.3212